OpenStudy (anonymous):

The speed of light is 1.86 × 105 miles per second. If the distance between the Sun and Earth is 9.3 × 107 miles, how long does it take sunlight to reach Earth? 2 seconds 0.002 seconds 500 seconds 50 seconds

OpenStudy (anonymous):

speed = distance/time \[1.86\times 10^5=\frac{ 9.3\times10^7 }{ t }\]